# ORCID: 0009-0007-8992-9682

### v2

@@ Abstract
Large language models have gained significant traction and popularity in recent times, extending their usage to code-generation tasks. While this field has garnered considerable attention, the exploration of testing and evaluating the robustness of code generation models remains an ongoing endeavor. Previous studies have primarily focused on code generation models specifically for the Python language, overlooking other widely-used programming languages. In this research, we conduct a comprehensive comparative analysis to assess the robustness performance of several prominent code generation models. Furthermore, we investigate how their performance varies across different programming languages. To accomplish this, we introduce perturbations in four key areas of the prompt: DocString, function name, syntax, and format. We have compiled and released a dedicated dataset for this purpose. This work presents our experimental findings, shedding light on the performance of code generation models in various scenarios.
Keywords:
Large Language Model Code Generation Robustness Testing
@@ 1 Introduction
Despite these advancements, the robustness of the LLMs on code generation tasks has not been extensively explored. LLMs are known to suffer from the issue of robustness in NLP tasks, i.e., studies(Samanta and Mehta, 2017; Gao et al., 2018; Li et al., 2018; Jin et al., 2020; Li et al., 2020) in NLP find that a slight change in the input can cause the model to generate a different prediction from the original result. Meanwhile, some recent studies  (Wang et al., 2023; Mastropaolo et al., 2023) focusing on LLM for code generation find that the Large Language Code generation model may also be sensitive to small perturbations on the prompt. For instance, Figure 1 demonstrates that even a minor function name modification can cause the model to produce incorrect code, although the non-perturbed prompt produced code that passed the test cases. Therefore, to test the robustness of code generation models, Wang et al. (2023) propose the first evaluation benchmark, ReCode, by customizing more than 30 semantic-preserved perturbations on the prompts (e.g., docstrings, function names), and then they evaluate the robustness of three code generation models (i.e., CodeGen, InCoder, and GPT-J).
However, the robustness evaluation (Wang et al., 2023) was limited to the Python language, raising the concern of the generalization to other programming languages, which is important to evaluate and show different facets (Austin et al., 2021), especially considering the diverse language features and different training resources across different programming languages. On the other hand, the evaluation benchmark they used is inadequate regarding the test cases, which may potentially impact the findings (Liu et al., 2023). For example, in Figure 2, the generated code in Java passes against the test suite in HumanEval-X but fails against the stronger test suite in EvalPlus (Liu et al., 2023).
@@
In this paper, we expand the current robustness assessment and revisit the findings from prior research to understand whether they still hold with different programming languages and under more rigorous testing. We aim to provide more insights to guide future research on improving the robustness of LLM-powered code generation. Specifically, we extend ReCode (Wang et al., 2022) from two aspects: First, in addition to Python, we evaluated the LLMs on more programming languages, including Java, C++, and JavaScript. Second, knowing that the test cases in HumanEval are inadequate, we construct an enhanced multi-lingual dataset of test cases based on EvalPlus (Liu et al., 2023). EvalPlus (Liu et al., 2023), an enhancement of the HumanEval (Chen et al., 2021) benchmark, introduces additional corner cases and test cases, leading to a more rigorous evaluation that expands a set of challenges, resulting in a higher failure rate for generated code by Large Language Models. Since EvalPlus test cases and solutions are in Python, we convert them to Java, C++, and JavaScript and manually update the solutions accordingly. The source code and benchmark dataset are publicly available. Our primary contributions are as follows.
@@
    We release the robustness evaluation datasets for Java, C++, and JavaScript languages, which are perturbed based on HumanEval-X, providing valuable resources that can be utilized for future research and related endeavors.
@@
    Our study reveals prevalent error types and vulnerabilities in code generation models when subjected to perturbations, highlighting the need for language-specific considerations and improvements in model design and evaluation.
@@
    We undertake the translation of the EvalPlus dataset from Python to Java, C++, and JavaScript. During this process, we also curate the test cases to ensure their compatibility with the mentioned programming languages. Additionally, we update the canonical solutions within the existing HumanEval-X datasets to align them with the translated test cases. The result of our efforts is a finalized dataset, encompassing translations and curated test cases, which we make publicly available.
In this study, we expand upon the work of ReCode by including Java, C++, and JavaScript languages, allowing for a comparative analysis. We assess the robustness of code generation models across these three languages, employing extended perturbations specific to each language. Four code generation models are experimented upon using nominal and perturbed datasets, evaluating their reactions and measuring robustness. We identify effective perturbation types and explore potential differences in model robustness based on programming languages. Additionally, we analyze prevalent error types resulting from the perturbations.
@@ 2.1 Background
In this part, we introduce two fundamental concepts essential to this work: (1) code generation, which is the task of generating source code based on the given natural language (NL) description and other information, (2) robustness testing, which aims to measure the ability of a system to function correctly despite the perturbation of the input.
Code generation aims to generate a code snippet based on a given prompt, which usually contains information about the method’s signature, an NL description of the code’s functionality, and test cases. The model is required to generate a complete code snippet based on the given information. Then the quality of the generated source code is evaluated against the unit tests. For example, in Figure 3, the prompt contains a docstring describing the method’s functionality and a method signature. The rest of the lines in that example are generated code by a code generation model.
@@
Robustness testing assesses the ability of a model to function correctly despite the perturbation of the input. During testing, a negligible perturbation is performed on the original prompt, and then the modified prompt is fed into the model to generate a new prediction, which is later compared with the original prediction.
In this paper, we mainly focus on the robustness of the code generation model, and during the testing, the prompt is modified slightly (e.g., by changing one letter from lowercase to uppercase), and then the model is asked to generate new code based on the new prompt. The correctness of the newly generated code is examined against the same test cases.
@@
Code generation models arrive from Large Language Models (Radford et al., 2019)(Brown et al., 2020) using few-shot learning for the code-generation tasks. These language models generally complete a sentence or verse. For code generation tasks, a prompt is pushed into a language model where the prompt has the information about a code’s parameters, test cases, signature, and other documentation, however, the implementation is missing. The language model completes the code by implementing it using the information from the given prompt.
@@ 2.2 Related Work
Nijkamp et al. (2022) introduces CodeGen, a series of transformer-based code generation models with various parameters for multiple languages, where the 16 billion parameter variant performs the best. CodeGeex by Zheng et al. (2023) is another multi-language-based series of models for code generation, where the largest model consists of 13 billion parameters. Other notable works in the field of code generation include Codex by Chen et al. (2021), Incoder by Fried et al. (2022), and GPT-J by Wang and Komatsuzaki (2021). Additionally, large language models such as GPT-3 by Brown et al. (2020) have shown potential for code-generation tasks. Furthermore, works such as CodeBERT by Feng et al. (2020) and PLBART by Ahmad et al. (2021) focus on training large language models specifically on code.
@@
Besides, recent studies have explored various aspects of code generation. Such as, Li et al. (2024a) and Li et al. (2024b) have focused on secure code generation, while Ling et al. (2024) investigate social bias in code generation. In addition, Lin et al. (2024) examine code generation through the lens of emulating the software development process, and Saha et al. (2024) propose methods for generating code using the source code from other programming languages.
@@ 3 Methodology
We applied semantic-preserving perturbations to two types of coding tasks: (1) code generation given a Docstring and method signature, and (2) code completion given a Docstring and an incomplete method with its first few lines. For code generation task, we include the perturbations to revise the docstrings and method signatures. and in the code completion task, from docstrings and incomplete methods. We applied two types of function and Docstring perturbations for the code generation task, and format and syntax perturbations for the code completion task.
@@ 3.1 Study Overview
This work focuses on the end-to-end code generation task under a black box setting, where it has access only to the model’s predictions with corresponding inputs. Figure LABEL:fig:overview illustrates the overall process of our study. We start with constructing the benchmark dataset for our experiments. Then, we apply 28 perturbation strategies from  Wang et al. (2022), focusing on four key areas in code. For code generation, we use perturbations on function names and docstrings, while for code completion, we apply them to format and syntax. These natural perturbations preserve the semantic meaning of the original prompts while reflecting common typographical errors developers may unintentionally introduce.
Finally, the original and the perturbed prompts are fed into the code generation model for evaluation separately. As we discussed earlier, the model takes a prompt as input and generates compilable code from the prompt. The prompt consists of a method signature, parameters, natural language comments, and others. In this work, we use an incomplete method as an input prompt, which consists of a docString and a method signature, which is identified as a nominal prompt throughout this work.
Note that this work does not cover adversarial attacks and only considers the robustness testing of the code generation models by generating perturbations randomly, providing them as input to the code generation models, and checking the performance with evaluation metrics.
@@ 3.2 Dataset Preparation
To address the limitations of Wang et al. (2022), which only evaluates code generation robustness in a single programming language (i.e., Python), we extend the evaluation to a multilingual setting by utilizing the HumanEval-X dataset. HumanEval-X, introduced by , serves as a benchmark for assessing the multilingual capabilities of code generation models.
@@
To improve evaluation robustness, we extend HumanEval-X by incorporating the enlarged test set provided by EvalPlus (Liu et al., 2023), which enhances HumanEval by refining NL descriptions, correcting solutions, and increasing the number of test cases per problem. Since EvalPlus is limited to Python, we converted its test cases into C++, Java, and JavaScript to facilitate multilingual evaluation. We also modified the canonical solutions in HumanEval-X to ensure compatibility with the additional test cases from EvalPlus, thereby enhancing the dataset’s comprehensiveness and robustness in evaluating code generation models.
@@
    We find it necessary to update certain canonical solutions, as they are unable to handle corner cases effectively. To address this, we adopt the approach outlined in EvalPlus to implement the necessary improvements in the solutions. After making these updates, we proceed to create the prompts using the revised versions of the solutions.
@@ 3.3 Semantic Preserved Perturbation
After preparing our dataset by converting the test cases to multiple programming languages and updating the canonical solutions accordingly, we apply semantically preserved perturbations to the prompts for each language in a distinct manner. For the code generation task, we perturb the prompt by applying eight function name perturbations and ten docstring perturbations. For the code completion task, we first construct incomplete code snippets using the prompt and a portion of the canonical solution. We then apply seven syntax perturbations and six formatting perturbations to these incomplete codes. The specific details of the perturbation techniques are described in the following sections.
@@ 3.3.1 Perturbation on DocString
Adaptation of Docstring Perturbations to multilingual To perturb a DocString, first, it is separated from the whole prompt, and the perturbation techniques are applied. The DocString may contain keywords of programming languages from code that must be untouched as changing them leads to altering the meaning. To maintain this, a blacklist of different programming languages is created which prevents the system from perturbing them. To separate the keywords and generate the blacklist, we utilize tree-sitter-java, tree-sitter-cpp, and tree-sitter-javascript for Java, C++, and JavaScript, respectively. The function names, variable names, and types are extracted from a code snippet by Tree-Sitter.
@@ 3.3.2 Perturbation on Function Name
With this scope of perturbation, the function name is changed slightly, alike the DocString. The six perturbations used for Function renaming are shown in Table 2 with examples and short explanations.
Among the six perturbations, InflectionalVariation and SynonymSubstituition are word-level, and the rest of them are character-level perturbations. CamelCase perturbation is applied to the C++ dataset as the function names in HumanEval-X for C++ are written following the snake case format. On the other hand, SnakeCase is applied to Java and JavaScript function names as they follow the camel case format in function names.

### v4

@@ Abstract
Large language models have gained significant traction and popularity in recent times, extending their usage to code-generation tasks. While this field has garnered considerable attention, the exploration of testing and evaluating the robustness of code generation models remains an ongoing endeavor. Previous studies have primarily focused on code generation models specifically for the Python language, overlooking other widely-used programming languages. In this work, we conduct a comprehensive comparative analysis to assess the robustness performance of several prominent code generation models and investigate whether robustness can be improved by repairing perturbed docstrings using an LLM. Furthermore, we investigate how their performance varies across different programming languages. To accomplish this, we introduce perturbations in four key areas of the prompt: DocString, function name, syntax, and format. We have compiled and released a dedicated dataset for this purpose. This work presents our experimental findings, shedding light on the performance of code generation models in various scenarios.
@@ 1 Introduction
Despite these advancements, the robustness of the LLMs on code generation tasks has not been extensively explored. LLMs are known to suffer from the issue of robustness in NLP tasks, i.e., studies(Samanta and Mehta, 2017; Gao et al., 2018; Li et al., 2018; Jin et al., 2020; Li et al., 2020) in NLP find that a slight change in the input can cause the model to generate a different prediction from the original result. Meanwhile, some recent studies  (Wang et al., 2023; Mastropaolo et al., 2023) focusing on LLM for code generation find that the Large Language Code generation model may also be sensitive to small perturbations on the prompt. For instance, Figure 1 demonstrates that even a minor function name modification can cause the model to produce incorrect code, although the non-perturbed prompt produced code that passed the test cases. Therefore, to test the robustness of code generation models, Wang et al. (2023) propose the first evaluation benchmark, ReCode, by customizing 28 semantic-preserved perturbations on the prompts (e.g., docstrings, function names), and then they evaluate the robustness of three code generation models (i.e., CodeGen, InCoder, and GPT-J).
Previous robustness evaluations of LLMs for code generation (Wang et al., 2023) have focused primarily on Python, leaving open the question of whether these findings generalize to other programming languages with distinct syntax, semantics, and uneven training resources (Austin et al., 2021). The benchmark used in prior work lacks sufficient test coverage, which can mask failures that appear under more rigorous evaluation suites (Liu et al., 2023), for instance, as shown in Figure 2, a Java solution may pass all HumanEval-X tests yet fail under the stronger EvalPlus suite. Another line of work has explored input smoothing techniques (Ye et al., 2020; Ji et al., 2024) to improve LLM performance on security-related instructions, but these methods are tailored to security-specific scenarios and do not address robustness issues arising from natural perturbations in code prompts. The broader problem of repairing or normalizing perturbed prompts, particularly perturbed docstrings to restore performance in code generation remains largely unexplored. This motivates our investigation into multilingual robustness, stronger evaluation settings, and post-perturbation prompt repair strategies for LLMs in code generation.
@@
In this paper, we expand the current robustness assessment and revisit the findings from prior research to understand whether they still hold with different programming languages and under more rigorous testing. We aim to provide more insights to guide future research on improving the robustness of LLM-powered code generation. Specifically, we extend ReCode (Wang et al., 2022) from two aspects: First, in addition to Python, we evaluated the LLMs on more programming languages, including Java, C++, and JavaScript. Second, knowing that the test cases in HumanEval are inadequate, we construct an enhanced multi-lingual dataset of test cases based on EvalPlus (Liu et al., 2023). EvalPlus (Liu et al., 2023), an enhancement of the HumanEval (Chen et al., 2021) benchmark, introduces additional corner cases and test cases, leading to a more rigorous evaluation that expands a set of challenges, resulting in a higher failure rate for generated code by LLMs. Since EvalPlus test cases and solutions are in Python, we convert them to Java, C++, and JavaScript and manually update the solutions accordingly. Besides this expanded robustness evaluation, we investigate whether robustness losses caused by prompt perturbations, particularly perturbed docstrings can be mitigated by automatically repairing the docstring using an LLM. The source code and benchmark dataset are publicly available.
Our primary contributions are as follows.
@@
    We significantly extend EvalPlus to EvalPlus-X to include a multi-language dataset. Particularly, we started with the HumanEval-X benchmark, and then we translated the EvalPlus test cases and modified the canonical solutions of HumanEval-X in Java, C++, and JavaScript to ensure correctness and compatibility;
  3. (3)
@@
    We release robustness evaluation datasets for Java, C++, and JavaScript, which are derived from our curated dataset based on HumanEval-X and EvalPlus test cases. These perturbed datasets provide valuable resources for future research and related endeavors.
@@
    We reveal prevalent linguistic and structural factors that drive robust drops in LLM code generation, showing how different programming languages and perturbation types amplify performance drops.
@@
    We experimented with an automatic mitigation strategy that uses LLMs to fix perturbed docstrings and demonstrate that it partially recovers robustness across models and languages.
In this study, we conduct a cross-language robustness evaluation of large language models for code generation across Java, C++, and JavaScript, aiming to provide a broader understanding of how LLMs behave when exposed to perturbed prompts in different programming languages. We evaluate six LLMs on nominal and perturbed datasets, measuring how robustness varies under different perturbation types and identifying language-dependent differences in performance.
We also introduce a mitigation approach that repairs perturbed docstrings using a secondary LLM prompt prior to code generation. This repair strategy is applied using Magicoder, chosen for its ability to interpret natural-language documentation. The results show slight robustness improvements for several perturbation types, with occasional degradation in others, revealing both the promise and the limitations of lightweight prompt-level repair techniques.
@@ 2.1 Background
In this part, we introduce two fundamental concepts essential to this work: (1) LLM code generation, which is the task of generating source code based on the given natural language (NL) description and other information, and (2) robustness testing on LLM code generation, which aims to measure the ability of a system to function correctly despite the perturbation of the input.
LLM code generation techniques leverage LLMs to generate a code snippet based on a given prompt, which usually contains information about the method’s signature, an NL description of the code’s functionality, and example test cases in the prompt. An LLM is required to generate a complete code snippet based on the given information. Often, the performance of LLM code generation is measured by benchmarks, such as HumanEval (Zheng et al., 2023), EvalPlus (Liu et al., 2023), etc. Such benchmarks utilize carefully crafted test cases to evaluate the generated code by LLMs. The example in Figure 3, shows cases that an LLM generates the code from the prompt that contains a docString describing the method’s functionality and a method signature.
@@
AI robustness is a known issue, and using LLM for code generation is no exception to the robustness challenge. In the context of LLM code generation, robustness testing is formed as techniques to assess an LLM’s performance degradation when a prompt is slightly perturbed while keeping the same semantics. During testing, a negligible perturbation is performed on the original prompt, and then the modified prompt is fed into the model to generate a new code snippet, which is later compared with the original prediction.
In this work, we focus on evaluating the robustness of code generation models. During robustness testing, we apply small, semantics-preserving modifications to the original prompt, for example, changing a single character in a function name, and then ask the model to generate code from this perturbed input. After generation, the correctness of the model’s output is assessed using the same test cases as in the nominal (unperturbed) setting. Following prior work, ReCode Wang et al. (2022), which studies robustness only in Python, we adapt and extend its 29 perturbation techniques to support C++, Java, and JavaScript.
@@
This work does not target adversarial manipulation of inputs. Instead, our focus is on robustness evaluation, where the goal is to examine whether code generation models can still perform reliably when nominal prompts are altered by natural or accidental variations. While adversarial attacks are designed to intentionally mislead models, the perturbations used in this study are semantically preserving and randomly generated, reflecting realistic user mistakes rather than crafted exploits. Our evaluation measures how model performance shifts under these natural perturbations using standard test pass metrics, allowing us to assess robustness degradation without adversarial intent.
@@ 2.2 Related Work
Nijkamp et al. (2022) introduces CodeGen, a series of transformer-based code generation models with various parameters for multiple languages, where the 16 billion parameter variant performs the best. CodeGeex by Zheng et al. (2023) is another multi-language-based series of models for code generation, where the largest model consists of 13 billion parameters. Other notable works in the field of code generation include Codex by Chen et al. (2021), Incoder by Fried et al. (2022), and GPT-J by Wang and Komatsuzaki (2021). Besides, LLMs such as GPT-3 by Brown et al. (2020) have shown potential for code-generation tasks. Furthermore, works such as CodeBERT by Feng et al. (2020) and PLBART by Ahmad et al. (2021) focus on training LLMs specifically on code.
@@
Besides, recent studies have explored various aspects of code generation. Such as, Li et al. (2024a) and Li et al. (2024b) have focused on secure code generation, while Ling et al. (2025) investigate social bias in code generation. Lin et al. (2024) examine code generation through the lens of emulating the software development process, and Saha et al. (2024) propose methods for generating code using the source code from other programming languages.
@@ 3 Methodology
We applied semantic-preserving perturbations to two types of coding tasks: (1) code generation given a Docstring and method signature, and (2) code completion given a Docstring and an incomplete method with its first few lines.
For code generation task, we include the perturbations to revise the docstrings and method signatures. and in the code completion task, from docstrings and incomplete methods. We applied two types of function and Docstring perturbations for the code generation task, and format and syntax perturbations for the code completion task.
@@ 3.1 Study Overview
This work focuses on the end-to-end code generation task under a black box setting, where it has access only to the model’s predictions with corresponding inputs. As depicted in Figure 4, the experimental pipeline proceeds through distinct stages following the prompt construction. First, the nominal and perturbed prompts for Java, C++, and JavaScript are processed by six distinct Large Language Models (LLMs): Incoder (1B and 6B), CodeGen-Multi (2B and 6B), Magicoder-S-DS-6.7B, and Qwen2.5-Coder-7B-Instruct.
The models generate code as output, which is then subjected to evaluation. To ensure a thorough assessment, the generated solutions are validated against both the original HumanEval test cases and the extended EvalPlus test cases. For RQ1, performance is quantified using three robustness metrics: Robust Pass (RPs@1), Robust Drop (RDs@1), and Robust Relative (RRs@1). These metrics are computed to produce a comparative report on model robustness.
For RQ2, we use 18 predefined features described in Section 5.2. These feature scores are computed from the generated code for both perturbed and nominal prompts, as well as from the prompts themselves. Using the pass/fail status and the feature values, we calculate impact scores for each language and each perturbation type. The results are then summarized to address the research question. Further details are provided in Section 5.2.
For RQ3, we focus on a single model, MagiCoder-S-DS-6.7B, selected from six evaluated models for its ability to repair docstrings. Only docstring perturbed datasets are considered. First, the model is prompted to repair the docstring within each perturbed prompt. The repaired docstring is then reintegrated into the prompt, and these repaired prompts are used to generate code following the same procedure as before. The generated code is then compared against code produced from the original perturbed docstrings. The results of this comparison are analyzed and visualized in RQ3.
Figure 4: Overview of the study
@@ 3.2 Dataset Preparation
To address the limitations of Wang et al. (2022), which only evaluates code generation robustness in a single programming language (i.e., Python), we extend the evaluation to a multilingual setting by utilizing the HumanEval-X dataset. HumanEval-X, introduced by Zheng et al. (2023), serves as a benchmark for assessing the multilingual capabilities of code generation models. The dataset includes Python, C++, Java, JavaScript, and Go. We exclude Go because it is less commonly used compared to the other languages in both research and industrial contexts. We also exclude Python from our evaluation since it was already thoroughly studied in prior work, ReCode Wang et al. (2022). We focused on C++, Java, and JavaScript, which represent diverse language paradigms (compiled vs. interpreted, statically vs. dynamically typed) and are widely used in real-world software development. This selection allows for a balanced and representative evaluation across different programming styles.
@@
To improve evaluation robustness, we extend EvalPlus to EvalPlus-X by incorporating the HumanEval-X prompts as well as the expanded test set from EvalPlus (Liu et al., 2023). EvalPlus enhances HumanEval by refining natural language descriptions, correcting canonical solutions, and increasing the number of test cases per problem. However, because EvalPlus is limited to Python, we translated its test cases into C++, Java, and JavaScript to enable multilingual evaluation. In addition, we revised the canonical solutions in HumanEval-X to ensure they remain consistent with the additional test cases introduced by EvalPlus. Together, these enhancements make EvalPlus-X a more comprehensive and robust benchmark for evaluating multilingual code generation models.
@@
    We find it necessary to update certain canonical solutions because they failed to handle corner cases, inputs at the edges of valid ranges or uncommon conditions that reveal latent bugs. Figures 2 and 8 show two representative failures: in Figure 2 the model produces a wrong answer for a corner input, while in Figure 8 the original solution leads to a timeout. Both examples demonstrate that passing the standard HumanEval-X tests does not guarantee robustness. To mitigate this, we follow the EvalPlus methodology to refine the affected canonical solutions manually by ourselves so they correctly handle these edge cases. After applying those fixes, we generate prompts from the revised solutions for subsequent evaluation.
@@ 3.3 Semantic Preserved Perturbation
After preparing our dataset by converting the test cases to multiple programming languages and updating the canonical solutions accordingly, we apply semantically preserved perturbations to the prompts for each language in a distinct manner. For the code generation task, we perturb the prompt by applying 7 function name perturbations and 10 docstring perturbations. For the code completion task, we first construct incomplete code snippets using the prompt and a portion of the canonical solution. We then apply 6 syntax perturbations and 6 format perturbations to these incomplete codes. The specific details of the perturbation techniques are described in the following sections.
@@ 3.3.1 Perturbation on DocString
Adaptation of Docstring Perturbations to multilingual To perturb a DocString, first, it is separated from the whole prompt, and the perturbation techniques are applied. The DocString may contain keywords of programming languages from code that must be untouched as changing them leads to altering the meaning. To maintain this, a blacklist of different programming languages is created which prevents the system from perturbing them. To separate the keywords and generate the blacklist, we utilize Tree-Sitter-Java, Tree-Sitter-C++, and Tree-Sitter-JavaScript for JavaSc, C++, and JavaScript, respectively. The function names, variable names, and types are extracted from a code snippet by Tree-Sitter.
@@ 3.3.2 Perturbation on Function Name
With this scope of perturbation, the function name is changed slightly, alike the DocString. The 7 perturbations used for Function renaming are shown in Table 2 with examples and short explanations. Among the 7 perturbations, the CamelCase perturbation is applied to C++ and JavaScript function names, where the original functions are written in snake case. Conversely, for Java, the snake case perturbation is applied since function names are written in camel case.
Among the 7 perturbations, InflectionalVariation and SynonymSubstituition are word-level, and the rest of them are character-level perturbations. CamelCase perturbation is applied to the C++ dataset as the function names in HumanEval-X for C++ are written following the snake case format. On the other hand, SnakeCase is applied to Java and JavaScript function names as they follow the camel case format in function names.